# Graph Residual based Method for Molecular Property Prediction

### v1

@@ 6.4 Results of GNN+VAE model
MicroAUC is favoured over MacroAUC in multilabel classification due to its ability to provide a balanced evaluation of model performance across all labels. Unlike MacroAUC, MicroAUC considers the contributions of each label equally, making it robust to class imbalance and sensitive to rare classes. By aggregating true positive and false positive rates, MicroAUC offers a comprehensive measure of the model’s discriminative ability, making it well-suited for assessing overall performance in multilabel classification tasks.
Apart from the AUCROC score, other major criteria include accuracy. Though the AUCROC is an important metric, it highlights only the model’s discriminatory power, it provides no information about the accuracy of the prediction.
In a multi-label setting, where instances can belong to multiple classes simultaneously, hamming accuracy $(H_{a})$ as described in Eq. 4 emerges as a more pertinent evaluation metric compared to traditional accuracy. Unlike accuracy, which evaluates the correctness of predictions on all labels collectively, Ha measures the fraction of correctly predicted labels per instance. $H_{a}$ accounts for the inherent complexity of multilabel classification tasks by assessing the model’s ability to correctly predict each label independently, offering a more comprehensive assessment of its predictive capabilities. $H_{a}$ is given by
@@
A representation of the Hamming Accuracy of the model on the test set has been plotted vs the epochs. This shows the model performance with epochs over the dataset. In Figure 13 both total accuracy scores and Hamming accuracy scores have been presented. The reason for the importance of Ha being more than the total accuracy is due to their formulation as described in Eq. 4. The total accuracy is stricter in the sense that it penalizes the entire set of labels of a data point in case of a mismatch in multi-label settings. Whereas, the HA considers each individual label in a label set as a single entity to be compared.
Figure 13: Plot of Hamming Accuracy vs Epochs to observe the improvement in accuracy of the model
Table 8: Predictions for Multi-label Multi-class Classification
@@
Apart from the normal prediction, we have had the idea of setting up separate labels for the missing values instead of directly imputing the values. This is highly effective for those datasets where the number of missing values is quite high in proportion to the other classes such as PCBA and ToxCast as can be seen from the statistical analysis in Figure 12 in Sec.6.2. The problem formulation now becomes a 3-class problem consisting of labels 0,1,2 which are then one-hot encoded in the vector space. The reason for adding a separate label is because some of these labels being missing may be inherently correlated with the property of the molecule, i.e. maybe those tests cannot be performed on that molecule to find the property.
@@
From the results, as shown in Table 8, a major increase in R2 values can be seen for Tox21 as well as ToxCast datasets. Furthermore, a substantial increase in total accuracy as well as Hamming accuracy can be seen. Since nothing specific has been provided on the kind of preprocessing that has been done in the papers [26], from the results thus it can be concluded that adding a separate class for the missing values seems more appropriate than imputing them.
7 Conclusions
@@
This work presents a new GRU-based Residual Graph model to overcome the limitations of the base ECC model. An end-to-end learning of the model has been presented on various SMILE-encoded molecular datasets, for both classification and regression tasks. From the Comparison of our model against the current SOTA approaches it has been observed that the model provides adequate improvement from the current SOTA, especially in cases of Lipophilicity and Boiling Point datasets. Apart from that, special focus has been placed in the context of multi-class multi-label classification of various datasets like SIDER, ToxCast and Tox21. Statistical analysis of the classification datasets including missing values and class imbalances have been presented. Both the methods of imputing missing values with 0 and adding them as a separate class have been tried. A quick comparison of the results shows that adding separate labels to the missing values helps improve the classification metrics. A VAE with the novel GNN as a backbone has been presented in this context, which has increased the accuracy of the association of labels with molecular feature embeddings. Our model performs quite competitively in this setting as well. The interpretability of our model and its relationship with the underlying chemistry has been shown here by the outputs of the GNNExplainer algorithm. Our model adheres better to the notion of the underlying chemistry than the base ECC model, as explained by the output graph due to the residual connections and GRU layers. A comprehensive study on the impact of residual connections, the number of layers, and its impact on property prediction will be undertaken as a future work.
Data availability
All the datasets that have been used have been mentioned and are publicly available. The compiled and formatted datasets can be provided upon valid requests.
Credit authorship contribution statement
@@
Abhishek Raj: Data Curation, Visualization, Validation, Investigation, Coding, Writing – original draft.
@@
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@ 6.4 Results of ECRGNN+VAE model
On distinct 85-5-10 train-validation-test splits, the ECRGNN demonstrates robust performance on sider, achieving a 76.7% hamming accuracy. Notably, the model excels on Tox21 with an impressive 60% accuracy and a consistent 92% hamming accuracy. The observed discrepancy is due to the lower label dimension in Tox21, contributing to the model’s optimal performance in this setting.
On the PCBA dataset, the ECRGNN underperforms with an AUC score of 75%, quite lower than the benchmark set 60 at 87%. The model showcases competitive performance in the expansive ToxCast dataset featuring an impressive 617 task labels, achieving a notable 78.2%
The versatility and adaptability of ECRGNN are evident in its consistent performance across these diverse datasets. The substantial enhancements in Micro-AUC-ROC scores, particularly in Sider, and the model’s ability to provide competitive results even for a high dimensional label space like that of ToxCast, underscore not only the efficacy of the approach but also its potential for accurate and comprehensive multilabel multiclass classification tasks in various domains. This, in turn, affirms the robustness and generalizability of the new GNN+VAE model. While ECRGNN excels at some tasks, some other datasets have better methods to deal with them individually. The upside to this is that ECRGNN performs consistently and competitively on all SMILE-encoded classification datasets with SOTA or near-SOTA performance.
MicroAUC is favoured over macroAUC in multilabel classification because it provides a balanced evaluation of model performance across all labels. Unlike MacroAUC, MicroAUC considers the contributions of each label equally, making it robust to class imbalance and sensitive to rare classes. By aggregating true positive and false positive rates, MicroAUC offers a comprehensive measure of the model’s discriminative ability, making it well-suited for assessing overall performance in multilabel classification tasks.
Apart from the AUCROC score, other major criteria include accuracy. Though the AUCROC is an important metric, it highlights only the model’s discriminatory power and provides no information about the accuracy of the prediction.
In a multi-label setting, where instances can belong to multiple classes simultaneously, hamming accuracy $(H_{a})$ as described in Eq. 4 emerges as a more pertinent evaluation metric compared to traditional accuracy. Unlike accuracy, which collectively evaluates the correctness of predictions on all labels, Ha measures the fraction of correctly predicted labels per instance. $H_{a}$ accounts for the inherent complexity of multilabel classification tasks by assessing the model’s ability to independently predict each label, offering a more comprehensive assessment of its predictive capabilities. $H_{a}$ is given by
@@
Figure 13: Plot of Hamming Accuracy vs Epochs to observe the improvement in accuracy of the model.
A representation of the Hamming Accuracy of the model on the test set has been plotted vs the epochs. This shows the model performance with epochs over the dataset. In Figure 13, both total accuracy scores and Hamming accuracy scores have been presented. The reason for the importance of Ha being more than the total accuracy is due to their formulation as described in Eq. 4. The total accuracy is stricter in the sense that it penalizes the entire set of labels of a data point in case of a mismatch in multi-label settings. Meanwhile, the HA considers each individual label in a label set as a single entity to be compared.
Apart from the normal prediction, the idea of setting up separate labels for the missing values instead of directly imputing the values is proposed. This is highly effective for those datasets where the number of missing values is quite high in proportion to the other classes, such as PCBA and ToxCast, as can be seen from the statistical analysis in Figure 12 in Sec.6.2. The problem formulation now becomes a 3-class problem consisting of labels 0, 1, and 2, which are then one-hot encoded in the vector space. The reason for adding a separate label is that some of these missing labels may be inherently correlated with the property of the molecule, i.e., maybe those tests cannot be performed on that molecule to find the property.
Table 8: Predictions for (Imputation 2) Classification.
@@
From the results, as shown in Table 8, a major increase in R2 values can be seen for Tox21 as well as ToxCast datasets. Furthermore, a substantial increase in total accuracy as well as Hamming accuracy can be seen. Since nothing specific has been provided on the kind of preprocessing that has been done in the papers 23, from the results, it can be concluded that adding a separate class for the missing values seems more appropriate than imputing them.
7 Discussion
To evaluate the effectiveness of ECRGNN, the results are benchmarked with the current state-of-the-art models: Elembert35, 32, GTOT23, BioAct-Het33, GMT34. Benchmarking was done for both classification and regression datasets.
7.1 Comparatative Study
Table 9: Comparison for Regression.
DATASET METRIC ECRGNN GMT+GCN GMT + ECRGNN Elembert GTOT
Boiling point RMSE 19.38 27.35 18.19 24.22 20.78
Melting Point RMSE 43.53 51.98 37.36 38.76 34.21
Lipo RMSE 0.1004 0.5411 0.1003 0.1162 0.1173
LogP RMSE 0.0334 0.1360 0.0789 0.1450 0.0926
Boiling point MAE 12.375 17.14 12.11 13.89 12.82
Melting Point MAE 30.915 42.031 26.09 25.65 26.17
Lipo MAE 0.0745 0.1192 0.0621 0.0880 0.079
LogP MAE 0.0256 0.0721 0.0249 0.0245 0.0251
Boiling point MAPE 2.72% 3.81% 2.65% 2.77% 2.74%
Melting Point MAPE 9.09% 11.55% 8.67% 8.45% 8.77%
Lipo MAPE 2.6% 4.2% 2.52% 2.9% 2.71%
LogP MAPE 6.99% 13.31% 6.91% 8.25% 8.27%
Table 9 shows a comparative study of regression experiments by various models to predict Boiling Point 4, Melting Point 4, Lipophilicity 4 and LogP 4. For accurate comparison, RMSE, MAE, and MAPE are presented. ECRGNN outperforms the other state-of-the-art models. If the pooling layer of the ECRGNN is replaced with GMT 34, the performance increases across many datasets. In the case of the Melting Point dataset 4, it has been noticed that Elembert 35 and GTOT 23 outperform the ECRGNN model. This may be because the Melting Point 4 has a much larger training set than others, and since transformers require a huge amount of data to work with, it follows that these model’s performance will improve with dataset size. However, it may be noted that ECRGNN can perform effectively even with small- and large-scale datasets.
Table 10: Comparison for Classification (0 imputation)
DATASET Metric ECRGNN GMT+GCN GMT+ECRGNN Elembert GTOT
Sider AUCROC 0.8321 0.8123 0.8412 0.778 0.6382
Tox21 AUCROC 0.7476 0.5389 0.66856 0.872 0.7236
ToxCast AUCROC 0.587 0.6021 0.5568 0.739 0.6782
Sider HA 0.7812 0.7491 0.75421 0.623 0.5146
Tox21 HA 0.92 0.891 0.926 0.918 0.9286
ToxCast HA 0.67 0.9621 0.9706 0.961 0.9768
Table 11: Comparison for Classification (2 imputation)
DATASET Metric ECRGNN GMT+GCN GMT+ECRGNN Elembert GTOT
Tox21 AUCROC 0.812 0.6023 0.6549 0.745 0.7613
ToxCast AUCROC 0.7491 0.6314 0.6012 0.681 0.7013
Tox21 HA 0.8806 0.7823 0.8706 0.798 -
ToxCast HA 0.8069 0.7694 0.9271 0.882 -
Table 10 and 11 provide comparative results of the approaches of imputing null values with 0 and 2, respectively. As can be seen, when the null values are imputed with 0, ECRGNN performs a bit poorly as compared to the state-of-the-art models due to the dataset being biased more towards 0. However, this factor is mitigated, and ECRGNN starts performing better with a separate classification category for the null values as presented in 11. Since the Sider 4 dataset does not have any missing values, ECRGNN outperforms the SOTA models in the corresponding dataset.
7.2 Limitations and Future Work
A detailed study of the novel ECRGNN, as well as the comparison, has been presented in this work. However, the study is not all-encompassing, and some limitations do exist. Though the GNN model has been fine-tuned, as seen from the regression results, the VAE architecture has not been hyper-parameter optimized for each corresponding dataset. Proper tuning of the hyperparameters is required for the VAE+GNN model to accurately measure the extent of ECRGNN’s capabilities. Proper ablation studies with various types of VAEs are also required.
The GRU-based ECRGNN model, as the highlight, improves the long-range molecular dependencies of the model. In this study, experimentation was done with only GRU-based mechanics. Although the GRU-based model provides improved performance, experimentation should also be done with other types of Recurrent Neural Networks. This has been reserved as a part of a future study. Apart from this, feature engineering and reduction methods to reduce the label space complexity are also among the directions intended to be pursued in a future study.
Conclusions
@@
This paper presents a new Residual Graph model, the ECRGNN, to overcome the limitations of the ECC and GNN models. An end-to-end learning and analysis of the model has been presented on various SMILE-encoded molecular datasets for both classification and regression tasks. ECRGNN provides adequate improvement from the current SOTA, especially in cases of the Lipophilicity and Boiling Point datasets. Apart from that, special focus has been placed on the context of multi-class, multi-label classification of various datasets like SIDER, ToxCast, and Tox21. Statistical analyses of used datasets and various imputation techniques have been compared, demonstrating that the addition of separate labels to the missing values helps improve the prediction results. A VAE with the ECRGNN as a backbone has been presented in this context, which has increased the accuracy of the association of labels with molecular feature embeddings. Ablation studies show the ECRGNN performs quite competitively in this setting as well. The relationship of the ECRGNN results with the underlying chemistry of the molecules has been shown through the GNNExplainer algorithm. ECRGNN is observed to adhere better to the notion of the underlying chemistry than the ECC model on which it is based. ECRGNN still lacks transparency and ablation studies from a variety of VAEs and other possible pipeline algorithms. A comprehensive study on the impact of residual connections, the number of layers, and its impact on property prediction will be undertaken as a future work.
Author contributions
@@
Abhishek Raj: Data Curation, Visualization, Validation, Investigation, Writing – original draft.
@@
Conflicts of interest
The authors declare no conflict of interest.
Data availability
All the datasets that have been used have been mentioned and are publicly available. The compiled and formatted datasets can be provided upon valid requests. {acknowledgement}
The authors thank Prof. Alankar Alankar for the guidance, support, and constructive feedback received during this project. The authors also thank IIT Bombay for providing us with the necessary computational resources.